# Toward Self-Improvement of LLMs via Imagination, Searching, and Criticizing

### v1

@@ 4.2 Data Synthesizing
Let ${\mathcal{D}}^{0}=\{({\bm{x}}_{i},{\bm{y}}_{i})\mid i\in[N]\}$ denote the initial dataset consisting of $N$ expert-generated prompt-response pairs. The data synthesizing process aims to expand this dataset by generating a set of synthesized prompts ${\mathcal{D}}^{1}=\{({\bm{x}}_{i}^{1},\cdots)\mid i\in[N]\}$. The generation of each synthesized prompt ${\bm{x}}_{i}^{1}$ can be mathematically described as a transformation $g$ applied to one or more examples from ${\mathcal{D}}^{0}$:
${\bm{x}}_{i}^{1}=g({\bm{x}}_{i_{1}}^{0},\cdots,{\bm{x}}_{i_{m}}^{0},\pi^{0})$
where ${\bm{x}}_{i_{1}}^{0},\cdots,{\bm{x}}_{i_{m}}^{0}$ are selected examples from ${\mathcal{D}}^{0}$. The transformation function $g$ controls the synthesis process, which can be a learnable function, manually defined heuristic rules, a strong LLM or the policy model itself $\pi^{0}$ equipped with data synthesis instructions. The data synthesizing process aims to enrich the diversity and complexity presented for the training of the policy model. Among various strategies, such as Self-instruct (Wang et al., 2022), Evol-instruct (Xu et al., 2023), we opt for a method akin to that described in Yu et al. (2023).
@@ 4.3.1 Option-level MCTS
When applying MCTS to LLMs, it is natural to perform token-level search , where each token is considered as an action (Liu et al., 2023). However, the substantial vocabulary size typical of LLMs presents a significant challenge i.e., conducting a deep search in such a vast space becomes increasingly complex as the search space expands exponentially. To mitigate this, some paper proposed a sentence-level search, treating each sentence or step as a search node (Feng et al., 2023). While this method reduces the search space, it might compromise the flexibility and effectiveness of applying MCTS to LLMs, which is particularly true for tasks where subtle variations in token can dramatically impact the outcome, or where a more comprehensive search beyond a sentence is necessary.
@@
Inspired by Sutton et al. (1999a); De Waard et al. (2016), we use the term option as a search node and propose option-level MCTS where each option represents a sequence of tokens, which can range from multiple tokens to several sentences. A comparisons of different levels search is listed in Table 1. Mathematically, an option $o=\langle{\mathcal{I}},\pi,\beta\rangle$, where ${\mathcal{I}}\subseteq{\mathcal{S}}$ is a set of initial states for the option; $\pi:{\mathcal{S}}\times{\mathcal{A}}\rightarrow[0,1]$ is a policy to generate actions, which in our case is a LLM; and $\beta:{\mathcal{S}}^{+}\rightarrow[0,1]$ is the termination function. Starting from a state $s_{t}$, we can choose all the options for which $s_{t}\in{\mathcal{I}}$. Once an option is chosen, the policy $\pi$ will generate actions for several steps until the option terminates according to the termination function $\beta$. As illustrated in Figure 2, option-level MCTS consists of the following operations:
  • Selection Starting from the root node, we iteratively select the child node based on Equation 1.
  • Simulation The scaled reward of the newly expanded node, as well as some simulated future trajectories are evaluated using the feedback functions, which will be discussed in §4.4.
Employing an option to substitute a single token within each node could reduces search space, as the number of options in a trajectory is much smaller than the number of tokens. This facilitates a deeper search, broader coverage of the search space, and minimizes the frequency of requesting feedback from functions such as the value model. Moreover, the option-level offers more flexibility compared to the sentence-level, as a new line can be treated as a special case of the termination function, as demonstrated in Table 1.
4.3.2 Importance Weighted Expansion
@@
In previous work related to option/sentence level tree search  Feng et al. (2023); Yao et al. (2024), it has been a common practice to assume that each node in the tree has the same predefined width i.e., branching factor. This is due to the fact that unlike token-level MCTS with a limited action space, the sample space at the option-level is exceedingly large, with an unlimited number of token combinations. Consequently, it is necessary to set a predefined maximum width. However, this assumption can often result in an inefficient search space, as it may be either too large or too small.
A more effective and efficient way to determine the branching factor for each node is to dynamically adjust it based on the importance of each node. This approach allows us to allocate a larger child budget to nodes of higher importance, thereby preventing inefficient exploration of these nodes and ensuring that we do not miss promising solutions. Meanwhile, by reducing the number of children for less important nodes, we can perform deeper searches at various levels of the tree, rather than considering all possible options at each node. Inspired by  Taylor et al. (2014); Clouse (1996), we define the importance of a node ${\bm{s}}_{t}$ as:
$I({\bm{s}}_{t})=\max_{{\bm{o}}_{t}}|v^{\pi}([{\bm{s}}_{t},{\bm{o}}_{t}])-v^{\pi}({\bm{s}}_{t})|$
@@
where $v^{\pi}$ is the value function which will be detailed in §4.4. $I({\bm{s}}_{t})$ captures the maximum value deviation from the current state. When this value is small, there is no need to explore further on this node, as there will not be a significant difference by rolling out on this node. Conversely, if the value is large, it is worth trying different children. We set the number of children allowed for a node $n({\bm{s}}_{t})$ to be linear with this importance, using a factor $\alpha$. In practice, to avoid extreme cases, we bound the number of children by depth-dependent constants $c_{\mathtt{min}}(t)$ and $c_{\mathtt{max}}(t)$:
$n({\bm{s}}_{t})=\max\left(c_{\mathtt{min}}(t),\min\left(\lfloor\alpha I({\bm{s}}_{t})\rfloor,c_{\mathtt{max}}(t)\right)\right).$
@@ 4.3.3 State Merge
With $n({\bm{s}}_{t})$ determined, another issue is that states under the same node can be very similar, causing many unnecessary sub-trees. To maximize diversity among states and cover as much space as possible with limited rollouts, we utilize the concept of move groups  Van Eyck & Müller (2012). By partitioning available options into distinct groups based on their similarities, with the maximum number of groups equal to the branching factor, we enhance diversity among groups. This strategy allows us to cover a larger problem space with limited search rollouts, making the search process more efficient.
In practice, each time we generate a new option from the policy, we use some heuristic functions to measure its similarity with existing options. The heuristic function can either be a faster rule-based measurement (e.g., edit distance) or a model-based method (e.g., prompting a LLM). Based on this, we decide whether to merge this option with a previous one or create a new group. This process is repeated until a maximum number of repetitions is reached. The details of this process are outlined in Algorithm 2.
Input max number of trails $max\_trials$, threshold $thres$
Output pool of children nodes
$n\leftarrow 0$
$min\_dist\leftarrow 0$
while $n<max\_trials$ and $min\_d\leq thres$ do
       ${\bm{o}}_{t}\sim\pi(s_{t})$
       $min\_d\leftarrow\min_{{\bm{o}}\in A_{t,\mathtt{pool}}}\mathtt{Dist}({\bm{o}}_{t},{\bm{o}})$
       $n\leftarrow n+1$
end while
Add ${\bm{s}}_{t+1}=[{\bm{s}}_{t},{\bm{o}}_{t}]$ to the pool of children nodes
Algorithm 2 Find Action with Minimum Distance Larger Than Threshold
In Algorithm 2, we iteratively sample an option ${\bm{o}}_{t}$ from the policy $\pi({\bm{s}}_{t})$ and compute the minimum distance $min\_d$ between ${\bm{o}}_{t}$ and the actions in the pool $A_{t,\mathtt{pool}}$ measured by distance function Dist. If $min\_d$ is larger than a predefined threshold $thres$ or the maximum number of trials $max\_trials$ is reached, the loop terminates and the resulting state ${\bm{s}}_{t+1}$ is added to the pool of children nodes.
@@ 4.3.4 Fast Rollout with Specialized LM
The simulation operation which employs a rollout policy to project future trajectories from a given state, is crucial for an effective MCTS. This process significantly improves the efficiency of exploration and exploitation, and enhances the accuracy of reward estimation [^1] Typically, the closer the simulation is to the termination state, the more accurate the reward estimation becomes.. By simulating numerous potential trajectories, MCTS can better approximate the likely outcomes of various actions, thereby facilitating a more informed and search process.
Ideally, $\pi_{\theta}$ would serve as the rollout policy, yet its computational demands render it impractical for the rapid simulations required by MCTS. To address this challenge, we propose the use of a smaller, specialized LM as the fast rollout policy $\pi^{\mathtt{fast}}$. Given a state ${\bm{s}}_{t}$, the fast rollout policy $\pi^{\mathtt{fast}}$ efficiently continues generation until it reaches a termination condition, denoted as $\pi^{\mathtt{fast}}({\bm{s}}_{t})$.
@@ 4.4 Critic
It is crucial for searching algorithms to have reliable guidance signals towards achieving the end goal. In AlphaLLM, we design three types of critic models to guide the search process, i.e., value function $v^{\pi}$ predicting the future reward, process reward models PRM estimating node quality, and outcome reward model ORM assessing the overall trajectory quality.
@@
The value function, denoted as $v^{\pi}({\bm{s}})$, is the expected reward starting from state ${\bm{s}}_{t}$ following the policy $\pi$ thereafter. To train a value function $v^{\pi}_{\phi}({\bm{s}})$ parameterized by $\phi$, we use the Monte Carlo (MC) estimate to empirically approximate the expected reward by averaging the rewards observed after many samplings starting from state $s$ and following policy $\pi$. The reward from a state is the sum of rewards obtained in the future, discounted by a factor $\gamma$ at each time step. Thus, the MC estimate of $v^{\pi}_{\phi}({\bm{s}})$ can be written as $v^{\pi}_{\phi}({\bm{s}})\approx\frac{1}{J}\sum_{j=1}^{J}G^{(j)}({\bm{s}})$ where $J$ is the number of trajectory starting from state ${\bm{s}}$, $G^{(j)}({\bm{s}})$ is the total discounted reward from state $s$ in the $j$-th trajectory. Particularly, given the expert demonstration dataset ${\mathcal{D}}=\{({\bm{x}}_{i},{\bm{y}}_{i})\}$, for each prompt ${\bm{x}}_{i}$, we generate trajectories ${\bm{\tau}}_{i}^{j}=\{{\bm{x}}_{i},{\bm{o}}_{i1}^{j},{\bm{o}}_{i2}^{j},\cdots,{\bm{o}}_{iT}^{j}\}$ by following policy $\pi$. A reward $r_{i}^{j}$ is assigned to indicate whether ${\bm{\tau}}_{i}^{j}$ aligns with ${\bm{y}}_{i}$, e.g., rewarding trajectories that contains correct answers in mathematical tasks or closely follows the instruction as the ground-truth. We then construct a dataset ${\mathcal{D}}_{\mathtt{value}}=\{({\bm{s}}_{it},v_{it})|i\in[N],t\in[T]\}$ in which ${\bm{s}}_{it}=[{\bm{x}}_{i}\cdot{\bm{o}}_{<it}]$ and $v_{it}=\frac{1}{J}\sum_{j=1}^{J}r^{j}_{iT}$. The value function $v_{\phi}^{\pi}$ is optimized by minimizing mean squared error:
${\mathcal{L}}_{\phi}=-{\mathbb{E}}_{({\bm{s}},v)\sim{\mathcal{D}}_{\mathtt{value}}}(v_{\phi}^{\pi}({\bm{s}})-v)^{2}$
We opt to initialize $v_{\phi}^{\pi}$ using the parameters from policy $\pi_{\theta}$, incorporating an MLP layer on top of it to output a scalar on each token. The scalar prediction at the last token of each state is used as the value.
@@
The value function often struggles with credit assignment problem (Sutton, 1984) and its learning could be inefficient due to delayed and sparse rewards (Sutton & Barto, 2018). Therefore, we propose to incorporate PRM that introduces process supervision (Lightman et al., 2023) for direct option assessment. PRM generates intrinsic rewards (Chentanez et al., 2004) to encourage explorations of advantageous options, effectively mitigating issues of reward sparsity by providing immediate, action-specific rewards. Given a state ${\bm{s}}_{t}$ and an option ${\bm{o}}_{t}$ at time $t$, the PRM aims to predict the immediate reward $r_{t}^{\texttt{PRM}}$ that results from taking option ${\bm{o}}_{t}$ in state ${\bm{s}}_{t}$. Formally, the PRM is a function $R({\bm{s}}_{t},{\bm{o}}_{t})\rightarrow r^{\mathtt{PRM}}_{t}$. Instead of adding a MLP layer on top of the policy model for outputting a scalar reward (Ouyang et al., 2022), we formulate PRM as a text generation task to best leverage LLM’s intrinsic knowledge for assessing the quality of an option. We use prefix sampling (Wang et al., 2023a) to estimate the quality of an option by starting from an option and exploring the final reward after reaching terminal states. The intuition is that an intermediate step can be regarded as a good if it frequently leads to achiving the goal. We adapt the dataset constructed for the value function as ${\mathcal{D}}_{\mathtt{PRM}}=\{({\bm{s}}_{it},{\bm{o}}_{t},r_{t}^{\mathtt{PRM}})|i\in[N],t\in[T]\}$ where $r_{t}^{\mathtt{PRM}}$ is the textual description of the reward, e.g., an option can be regarded as good if $v_{it}$ is larger than certain threshold. To train PRM, we initialize it from the policy model $\pi$ and use the following prompt templates and typical language model loss.

### v2

@@ 4.2 Data Synthesizing
Let ${\mathcal{D}}^{0}=\{({\bm{x}}_{i},{\bm{y}}_{i})\mid i\in[N]\}$ denote the initial dataset consisting of $N$ expert-generated prompt-response pairs. The data synthesizing process aims to expand this dataset by generating a set of synthesized prompts ${\mathcal{D}}^{1}=\{({\bm{x}}_{i}^{1},\cdots)\mid i\in[N]\}$. The generation of each synthesized prompt ${\bm{x}}_{i}^{1}$ can be mathematically described as a transformation $g$ applied to one or more examples from ${\mathcal{D}}^{0}$, ${\bm{x}}_{i}^{1}=g({\bm{x}}_{i_{1}}^{0},\cdots,{\bm{x}}_{i_{m}}^{0},\pi^{0})$ where ${\bm{x}}_{i_{1}}^{0},\cdots,{\bm{x}}_{i_{m}}^{0}$ are selected examples from ${\mathcal{D}}^{0}$. The transformation function $g$ controls the synthesis process, which can be a learnable function, manually defined heuristic rules, a strong LLM or the policy model itself $\pi^{0}$ equipped with data synthesis instructions. The data synthesizing process aims to enrich the diversity and complexity presented for the training of the policy model. Among various strategies, such as Self-instruct (Wang et al., 2022), Evol-instruct (Xu et al., 2023), we opt for a method akin to that described in Yu et al. (2023).
@@ 4.3.1 Option-level MCTS
When applying MCTS to LLMs, it is natural to perform token-level search, where each token is considered as an action (Liu et al., 2023). However, the substantial vocabulary size typical of LLMs presents a significant challenge i.e., conducting a deep search in such a vast space becomes increasingly complex as the search space expands exponentially. To mitigate this, some efforts proposed a sentence-level search, treating each sentence or step as a search node (Feng et al., 2023). While this method reduces the search space, it might compromise the flexibility and effectiveness of applying MCTS to LLMs, which is particularly true for tasks where subtle variations in token can dramatically impact the outcome, or where a more comprehensive search beyond a sentence is necessary.
@@
Inspired by Sutton et al. (1999a); De Waard et al. (2016), we use the term option as a search node and propose option-level MCTS where each option represents a sequence of tokens, which can range from multiple tokens to several sentences. A comparisons of different levels search is listed in Table 1. Mathematically, an option $o=\langle{\mathcal{I}},\pi,\beta\rangle$, where ${\mathcal{I}}\subseteq{\mathcal{S}}$ is a set of initial states for the option; $\pi:{\mathcal{S}}\times{\mathcal{A}}\rightarrow[0,1]$ is a policy to generate actions, which in our case is a LLM; and $\beta:{\mathcal{S}}^{+}\rightarrow[0,1]$ is the termination function. Starting from a state $s_{t}$, we can choose all the options for which $s_{t}\in{\mathcal{I}}$. Once an option is chosen, the policy $\pi$ will generate actions for several steps until the option terminates according to the termination function $\beta$. The option-level MCTS consists of stages including selection, expansion, simulation, and backpropagation. The option-level formulation offers more flexibility compared to the sentence-level, as a new line can be treated as a special case of the termination function, as demonstrated in Table 1. Additional detailed steps of the option-level MCTS can be found in Appendix A.2.
4.3.2 Importance-Based Adaptive Branching
@@
In previous works related to option/sentence level tree search  (Feng et al., 2023; Yao et al., 2024), it was a common practice to assume that each node in the tree has the same predefined width, i.e., branching factor. This assumption was due to the fact that unlike token-level MCTS with a limited action space, the sample space at the option-level is exceedingly large, with an unlimited number of token combinations. As a result, it was necessary to set a predefined maximum width for each node. However, this predefined branching factor is hard to set, as an improper choice can lead to a search tree that is either too shallow or too thin, resulting in an inefficient exploration of the search space.
To quantify the error induced by the branching factor limit, we defined the branching error $E_{\phi}(t)$. For a node $t$ with a branching factor of $m_{t}$, it aims to use the $m_{t}$ child options ${\bm{o}}_{t}^{i}\sim{\mathcal{D}}_{t}^{children}$ (where $i\in\{1,\ldots,m_{t}\}$) to represent all possible options. Consequently, for a legal option ${\bm{o}}_{t}^{j}\sim\pi({\bm{s}}_{t})$ from the option space, we can calculate the minimal value difference between it and the $m_{t}$ existing options, which captures the error associated with representing other possible options using the $m_{t}$ available options. It can be formulated as $E_{\phi}(t)=\mathop{\mathbb{E}_{{\bm{o}}_{t}^{j}\sim\pi({\bm{s}}_{t})}}[\min_{{\bm{o}}_{t}^{i}}|v_{\phi}^{\pi}([{\bm{s}}_{t},{\bm{o}}_{t}^{j}])-v_{\phi}^{\pi}([{\bm{s}}_{t},{\bm{o}}_{t}^{i}])|]$, where $v_{\phi}^{\pi}$ is the value function which will be detailed in §4.4. Here we define the importance of node ${\bm{s}}_{t}$ as $I({\bm{s}}_{t})=\max_{{\bm{o}}_{t}^{i}}|v_{\phi}^{\pi}([{\bm{s}}_{t},{\bm{o}}_{t}^{i}])-v_{\phi}^{\pi}({\bm{s}}_{t})|.$ For simplicity, we assume that the value of the children nodes are uniformly distributed (a detailed analysis of the Gaussian distribution can be found in Appendix A.4). Under this assumption, we show in Appendix A.3 that $E_{\phi}(t)\leq\frac{I({\bm{s}}_{t})}{m_{t}-1}.$ While $E_{\phi}$ is less than some $\epsilon$, we aim to use a smaller total number of nodes for efficiency.
Theorem 4.1.
The optimal branching factor $m_{t}$ in a tree search is set such that $m_{t}-1$ is proportional to the node importance $I({\bm{s}}_{t})$, under the condition $\frac{I({\bm{s}}_{t})}{m_{t}-1}\leq\epsilon$. Refer to Appendix A.3 for the detailed proof.
@@
A similar concept has also been proposed in  Taylor et al. (2014); Clouse (1996). Intuitively, $I({\bm{s}}_{t})$ captures the maximum value deviation from the current state. When this value is small, there is no need to explore further on this node, as there will not be a significant difference by rolling out on this node. Conversely, if the value is large, it is worth trying different children. We set the number of children allowed for a node $n({\bm{s}}_{t})$ (after extracting $1$) to be linear with this importance, using a factor $\alpha$. In practice, to avoid extreme cases of large variance of $I({\bm{s}}_{t})$ in the early stage, we bound the number of children by depth-dependent constants $c_{\mathtt{min}}(t)$ and $c_{\mathtt{max}}(t)$, $n({\bm{s}}_{t})=\max\left(c_{\mathtt{min}}(t),\min\left(\lfloor\alpha I({\bm{s}}_{t})\rfloor+1,c_{\mathtt{max}}(t)\right)\right).$
@@ 4.3.3 State Merge
With $n({\bm{s}}_{t})$ determined, another issue is that options under the same node may be very similar, leading to many unnecessary sub-trees. Since we cannot directly control the ${\bm{o}}_{t}\sim\pi({\bm{s}}_{t})$, one strategy to mitigate this issue is to utilize the concept of move groups, as discussed in  Van Eyck & Müller (2012). By merging similar nodes into the same group, we can increase the diversity among groups, thereby covering a larger problem space with limited search rollouts and making the search process more efficient.
Here, we adapt the definition of node predicate $p_{vM}$ from  Abel et al. (2018) and  Fu et al. (2024) to represent whether two nodes are extremely similar. In practice, each time we generate a new option from the policy, we use heuristic functions as $p_{vM}$ to check its similarity with all existing groups. The heuristic function can either be a faster rule-based measurement (e.g., edit distance) or a model-based method (e.g., prompting a language model). Based on this, we decide whether to merge this option with a previous one or create a new group.
@@ 4.3.4 Fast Rollout with Specialized LM
The simulation operation which employs a rollout policy to project future trajectories from a given state, is crucial for an effective MCTS. This process significantly improves the efficiency of exploration and exploitation, and enhances the accuracy of reward estimation [^1] Typically, the closer the simulation is to the termination state, the more accurate the reward estimation becomes.. Estimations made at the end of trajectories tend to have lower bias but higher variance; thus, simulating multiple possible trajectories yields low-bias, low-variance estimates, enabling a more informed and effective search process. Ideally, $\pi_{\theta}$ would serve as the rollout policy, yet its computational demands render it impractical for the rapid simulations required by MCTS. To address this challenge, we propose the use of a smaller, specialized LM as the fast rollout policy $\pi^{\mathtt{fast}}$. Given a state ${\bm{s}}_{t}$, the fast rollout policy $\pi^{\mathtt{fast}}$ efficiently continues generation until it reaches a termination condition, denoted as $\pi^{\mathtt{fast}}({\bm{s}}_{t})$.
@@ 4.4 Critic
In AlphaLLM, we design three types of critic models to guide the search process.
@@
The value function, denoted as $v^{\pi}({\bm{s}})$, represents the expected return starting from state ${\bm{s}}$ and following policy $\pi$ thereafter, given by $v^{\pi}({\bm{s}})=\mathop{\mathbb{E}}_{\tau\sim\pi}[R(\tau)|s_{0}={\bm{s}}]$ where $R(\tau)$ represents the discounted return of trajectory $\tau$. To train a parameterized value function $v^{\pi}_{\phi}({\bm{s}})$, given the prompts ${\mathcal{D}}=\{({\bm{x}}_{i},\cdots)\mid i\in[N]\}$, for each prompt ${\bm{x}}_{i}$, we generate multiple trajectories ${\bm{\tau}}_{i}^{j}=\{{\bm{x}}_{i},{\bm{o}}_{i1}^{j},{\bm{o}}_{i2}^{j},\cdots,{\bm{o}}_{iT}^{j}\}$ by following policy $\pi$ for $J$ times. A final reward $r_{i}^{j}$ is assigned to indicate whether ${\bm{\tau}}_{i}^{j}$ aligns with ${\bm{y}}_{i}$—for example, rewarding trajectories that contain correct answers in mathematical tasks or closely follow instructions as ground truth. We then construct a dataset ${\mathcal{D}}_{\mathtt{value}}=\{({\bm{s}}^{j}_{it},v^{j}_{it})\mid i\in[N],t\in[T],j\in[J]\}$ where ${\bm{s}}^{j}_{it}=[{\bm{x}}_{i}\cdot{\bm{o}}^{j}_{<it}]$ and $v^{j}_{it}=r^{j}_{i}$. The value function $v_{\phi}^{\pi}$ is optimized by minimizing the mean squared error: ${\mathcal{L}}_{\phi}=-{\mathbb{E}}_{({\bm{s}},v)\sim{\mathcal{D}}_{\mathtt{value}}}(v_{\phi}^{\pi}({\bm{s}})-v)^{2}$. Similar to  (Feng et al., 2023), $v_{\phi}^{\pi}$ is a LLM with an MLP layer on top to output a scalar on each token, using the scalar prediction at the last token of each state as the value.
@@
The value function often struggles with credit assignment problem (Sutton, 1984) and its learning could be inefficient due to delayed and sparse rewards (Sutton & Barto, 2018). Therefore, we propose to incorporate PRM that introduces process supervision (Lightman et al., 2023) for direct option assessment. PRM generates intrinsic rewards (Chentanez et al., 2004) to encourage explorations of advantageous options, effectively mitigating issues of reward sparsity by providing immediate, action-specific rewards. Given a state ${\bm{s}}_{t}$ and an option ${\bm{o}}_{t}$ at time $t$, the PRM aims to predict the immediate reward $r_{t}^{\texttt{PRM}}$ that results from taking option ${\bm{o}}_{t}$ in state ${\bm{s}}_{t}$. Formally, the PRM is a function $R({\bm{s}}_{t},{\bm{o}}_{t})\rightarrow r^{\mathtt{PRM}}_{t}$. While PRM ideally requires quality labels for each state  (Uesato et al., 2022), due to the high cost and time involved in obtaining these, MC estimation with prefix sampling (Wang et al., 2023) is used as a proxy, which aligns with the objective of the value function. Instead of adding a MLP layer on top of the policy model for outputting a scalar reward (Ouyang et al., 2022), we formulate PRM as a text generation task to best leverage LLM’s intrinsic knowledge for assessing the quality of an option. We adapt the dataset constructed for the value function as ${\mathcal{D}}_{\mathtt{PRM}}=\{({\bm{s}}_{it},{\bm{o}}_{t},r_{t}^{\mathtt{PRM}})|i\in[N],t\in[T]\}$ where $r_{t}^{\mathtt{PRM}}$ is the textual description of the reward, e.g., an option can be regarded as good if $v_{it}$ is larger than certain threshold. To train PRM, we initialize it from the policy model $\pi$ and use the following prompt templates and typical language model loss. The prompt template is shown in Appendix A.5.
ORM
In additional to the value function and PRM, ORM is also used to guide MCTS. ORM is designed to evaluate options sequences in their entirety, assessing the extent to which the complete trajectory aligns with the desired end goal (Uesato et al., 2022; Lightman et al., 2023; Wang et al., 2023; Feng et al., 2023). The outcome evaluation complements value function and PRM by offering a comprehensive assessment of trajectories. Crucially, ORM plays a vital role in the simulation stage of MCTS by providing more accurate signals on the terminal state, which in turn facilitates a more balance between exploration and exploitation strategies. ORM is formulated as a text generation task, similar to PRM. We leverage the same dataset for the value function training and construct ${\mathcal{D}}_{\mathtt{ORM}}=\{({\bm{x}}_{i},{\bm{o}}_{1:T}^{i},r_{i}^{\mathtt{ORM}})|i\in[N]\}$, where each instance includes a initial state or prompt ${\bm{x}}_{i}$, a sequence of actions or options ${\bm{o}}_{1:T}^{i}$ taken from that state, and a textual reward $r_{i}^{\mathtt{ORM}}$ indicating the sequence’s success or quality. Similarly, ORM is initialized from the policy model $\pi$ and the following prompt templates and language model loss are used for training. The prompt template is shown in Appendix A.5.
The final score evaluation of a state ${\bm{s}}$ is a weighted sum of the value function, PRM, and ORM: $s({\bm{s}})=\beta_{\text{value}}\cdot v_{\phi}^{\pi}({\bm{s}})+\beta_{\text{PRM}}\cdot\texttt{PRM}{}({\bm{s}})+\beta_{\text{ORM}}\cdot\mathbb{E}_{\tau\sim\pi^{\mathtt{fast}}({\bm{s}})}[\texttt{ORM}{}(\tau)]$, where $\tau\sim\pi^{\mathtt{fast}}({\bm{s}})$ represents trajectories starting from ${\bm{s}}$ under $\pi^{\mathtt{fast}}$, and $\beta_{\text{value}}$, $\beta_{\text{PRM}}$, $\beta_{\text{ORM}}$ are hyperparameters. In practice, we found that the value function model has better precision and calibration, while PRM has superior recall (Appendix A.10). Although ORM with fast rollouts provides low-bias, low-variance estimates, it still inherits some bias from $\pi^{\mathtt{fast}}$. Thus, combining these critics yields a stronger evaluation signal.
4.5 Policy Self-Improvement
The policy improvement an iterative process with each iteration containing two main steps: data generation and policy finetuning.
Data generation
In this step, we assume to have the current policy $\pi_{\theta_{k}}$ and synthetic prompts ${\mathcal{D}}_{k}=\{{\bm{x}}^{k}_{1},\dots\}$ at the $k$-th round, where each ${\bm{x}}^{k}_{1}$ represents a question. We obtain the corresponding training data ${\mathcal{D}}_{k}$ for policy $\pi_{\theta_{k}}$ by firstly performing $\eta$Mcts on ${\mathcal{D}}_{k}$4.3) and then sampling a trajectory ${\bm{y}}^{k}_{i}$ from the corresponding tree for each question ${\bm{x}}^{k}_{i}$. Here we choose the trajectory that yield the highest critic score on the leaf node for each input question. Next, we filter out instances where the corresponding trajectory is substandard forming ${\mathcal{D}}_{k}=\{({\bm{x}}^{k}_{i},{\bm{y}}^{k}_{i})~{}|~{}f({\bm{x}}^{k}_{i},{\bm{y}}^{k}_{i})>\gamma\}$ where $f$ represents a function for quality scoring, and $\gamma$ indicates a threshold. There can be several ways to implement the function, and here we simply use the ORM4.4).
Policy finetuning
With the obtained training data ${\mathcal{D}}_{k}$, we organize the data into the prompt templates shown in Appendix A.5. Then the policy $\pi_{\theta_{k}}$ is finetuned using target-loss: $\mathcal{L}_{\theta_{k}}=\mathbb{E}_{({\bm{x}}^{k}_{i},{\bm{y}}^{k}_{i})\sim{\mathcal{D}}_{k}}\big{[}\log\pi_{\theta_{k}}({\bm{y}}^{k}_{i}|{\bm{x}}^{k}_{i})\big{]}$, resulting in an updated policy $\pi_{\theta_{k+1}}$. We leave other training methods, such as DPO (Rafailov et al., 2023) or PPO (Schulman et al., 2017) in future work.
5 Experiments
5.1 Experiment Setups
AlphaLLM is generally applicable to a wide spectrum tasks. As an early exploration, in this paper, we conduct experiments on mathematical reasoning problems where the learning signals are clear to define i.e., , final answer is correct or wrong. We choose to evaluate on two widely used datasets GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). For GSM8K, we utilize the whole test set while for MATH, due to computation constraints, we utilize a subset following the same procedure of Lightman et al. (2023). We evaluate the performance of predicting answers correctly for policy models. In addition, we calculate the average rollouts, represented by the number of nodes in the tree, as a measure of computational efficiency. We compare the performance of AlphaLLM with a suite of proprietary model, including OpenAI’s GPT-4 and GPT-3.5, Anthropic’s Claude-2, as well as Google’s PaLM-2 and the gemini model family. To ensure a fair and consistent evaluation, we employ CoT as our primary prompting method. Additionally, we conduct comparisons with strong open-source models, including Llama-2-70b (Touvron et al., 2023a) and WizardMath-70B-V1.0 (Luo et al., 2023).
We select Llama-2-70b as the policy model for the GSM8K dataset and WizardMath-70B-V1.0 for the MATH dataset. To construct the training dataset for the value function, PRM and ORM, we generate 50 trajectories for each prompt and construct the training target following Section 4.4. Both PRM and ORM are initialized using the weights from the policy model, while the value function uses a smaller Llama-2-13b model, as we observed no performance gains from increasing the value function model size. In the design of ORM, tool usage is not incorporated for GSM8K. However, for MATH, we enhance ORM by incorporating tools like python sympy to assess the quality of a trajectory, in a manner similar to that described by Gou et al. (2023). The training employ a learning rate of 1e-6 and are trained for one epoch. For the fast rollout policy model, we opt for the Abel-002-7B model (Chern et al., 2023) for both the GSM8K and MATH tasks for its high efficiency and superior performance. For the MCTS parameters, they are configured at different scales, as shown in Appendix A.6. We set $\beta_{\text{value}}$, $\beta_{\text{PRM}}$, and $\beta_{\text{ORM}}$ all to 1.0.
For policy self-improving (§4.5), we train the policy model up to 3 epochs, setting batch size to 128, learning rate to $5\times 10^{-6}$ and minimal learning rate to $1\times 10^{-6}$. Linear warm-up and decay is used with warm-up percent to be 10%. We perform early stopping based on a devset held out from the training instances. For GSM8K experiments, we perform two rounds of self-improving, synthesizing 6.4k and 7.9k prompts(Yu et al., 2023) respectively to obtain the corresponding MCTS outputs for training. For MATH experiments, we only perform one round of self-improving due to limited computation resources, and 5.9k prompts are synthesized.